OpenStudy (anonymous):

Find the unknown number. Three added to 4 times a number is equal to 5 less than 6 times the number. (Points : 1) –8 –4 4 8

OpenStudy (anonymous):

3 + 4x = 6x - 5

OpenStudy (secret-ninja):

One easy way to figure this out is to plug in the choices. Lets try -8, shall we? 3 = 4 = 7. 7 x -8 = -1. The nest part is kind of tricky. You need to see if -1 is equal to 5 minus 6, but you need to multiply 6 by the number. 6 x -8 = -48. -1 is not 5 less than -48. This means your answer will not be A. Do you understand, or do you need me to stay?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Answer is C.

OpenStudy (anonymous):

i think i understand??

OpenStudy (anonymous):

look at the first comment that is the correct formula, and reread the question

OpenStudy (secret-ninja):

Thanks a lot @RyleeC It say specifically in our Code of Conduct- " Give Help, not Answers. I will encourage and guide those needing help, and not just give them an answer." http://openstudy.com/code-of-conduct You have to help them through it or they will never learn!

OpenStudy (anonymous):

That's why I gave them the formula...

OpenStudy (secret-ninja):

If Rylee think that is the answer, lets try it out.

OpenStudy (anonymous):

Chill out. Its correct.

OpenStudy (secret-ninja):

3 + 4 = 7. 7 x 4 = 28. 6 x 4= 24 28 - 5= 23 It equaled 23, not 24.

OpenStudy (secret-ninja):

@RyleeC

OpenStudy (anonymous):

Lol your wrong again bud.

OpenStudy (anonymous):

3 + 4x = 6x - 5

OpenStudy (anonymous):

19= 6x -5

OpenStudy (anonymous):

19 = 19

OpenStudy (secret-ninja):

I apologize, I found my mistake. Good job Rylee. Just remember to not even give out the formula, help them form it themselves. :)

OpenStudy (anonymous):

You shouldn't help others if you don't even understand the math yourself. So next time you want to tell me how to teach, check your own work.

OpenStudy (secret-ninja):

I was trying to help, and yes I did understand. I was trying to help you both, found my mistake, and admitted to it. There is no need to be rude when I have done nothing wrong. Is it my patience that you find infuriating?

OpenStudy (secret-ninja):

@RyleeC You should answer when someone is speaking to you.

OpenStudy (anonymous):

My computer is slow, so it takes awhile to load. Firstly most people who post question have online classes, and test and quizzes that are timed. I give the answer and explain later. They need help fast, and have infinite time to learn after.

OpenStudy (anonymous):

I meant no disrespect towards you, but it is your responses towards me that has lost any respect that I had of you.

OpenStudy (secret-ninja):

If you do that, it will only make them want to get the answer, and (almost all of the time) leave when they have it. I have a feeling this conversation could go on for a long time, so lets just move on. Sorry if I caused you any anger or discomfort. Have a nice day.